A father is now three times as old as his son. Twelve years ago he was six times as old as his son. Find the Present age of each
Found 2 solutions by algebrahouse.com, ewatrrr:
Answer by algebrahouse.com(1659)

"A father is now three times as old as his son. Twelve years ago he was six times as old as his son. Find the Present age of each"

x = son's age now
3x = father's age now {father is 3 times as old as his son}

x - 12 = son's age twelve years ago
3x - 12 = father's age twelve years ago

3x - 12 = 6(x - 12) {twelve years ago, father was six times as old as his son}
3x - 12 = 6x - 72 {used distributive property}
3x = 6x - 60 {added 12 to both sides}
-3x = -60 {subtracted 6x from both sides}
x = 20 {divided both sides by -3}
3x = 60 {substituted 20, in for x, into 3x}

son is 20 now
father is 60 now
www.algebrahouse.com

Answer by ewatrrr(24785)


Hi
A father is now three times as old as his son
Let x and 3x represent the ages of the son and father NOW repectively
Question states*** 12years ago...
3x-12 = 6(x-12)
Solving for x
3x-12 = 6x - 72
60 = 3x
20 = x, son's age now. The father is now 60.
CHECKING our Answer***12years ago...
48 = 6*8